# Unleashing the Power of LLMs in Dense Retrieval with Query Likelihood Modeling

### v1

@@ Abstract.
submissionid: 517copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Language modelsccs: Information systems Learning to rankccs: Information systems Novelty in information retrieval
@@ 5.2. Training Settings
Table 2. MS MARCO-dev passage retrieval (performance measured by MRR@10, Recall@1000, NDCG@10). means that we reproduce the models using LlaMa2 according to the original paper. ‘-’ indicates that the original paper does not specifically mention the corresponding terms. ‘’ indicate significant improvements over Summarize, Echo, LLM2VEC (p¡0.05) and Hommel’s correction ($\alpha$ = 0.05).
@@
RepLLaMA (FT) (Ma et al., 2024) No Static(CoCondenser+BM25) 7B hard 41.6 - 72.8 69.9
@@
LLM2VEC (LoRA+LoRA) (BehnamGhader et al., 2024) No Static(CoCondenser+BM25) 7B hard
LLM2VEC (FT+LoRA) (BehnamGhader et al., 2024) No Static(CoCondenser+BM25) 7B hard 41.9 99.3 72.9 74.4
@@
LLM-QL (FT+LoRA) No Static(CoCondenser+BM25) 7B hard 42.4 99.4 73.6 72.4
@@ 7. Further Analyses
In this section, we conduct a thorough analysis of LLM-QL to clarify its advantages.

### v2

@@ Abstract.
submissionid: 5376copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Language modelsccs: Information systems Learning to rankccs: Information systems Novelty in information retrieval
@@ 5.2. Training Settings
Table 2. MS MARCO-dev passage retrieval (performance measured by MRR@10, Recall@1000, NDCG@10). means that we reproduce the models using LlaMa2 according to the original paper. ‘-’ indicates that the original paper does not specifically mention the corresponding terms. ‘’ indicate significant improvements over Summarize, Echo, LLM2VEC (p¡0.05) and Hommel’s correction ($\alpha$ = 0.05). We used full-parameter fine-tuning (Full) and LoRA fine-tuning (LoRA) strategies, respectively.
@@
RepLLaMA (Full) (Ma et al., 2024) No Static(CoCondenser+BM25) 7B hard 41.6 - 72.8 69.9
@@
LLM2VEC (LoRA+LoRA) (BehnamGhader et al., 2024) No Static(CoCondenser+BM25) 7B hard 41.8 99.3 74.4 72.5
LLM2VEC (Full+LoRA) (BehnamGhader et al., 2024) No Static(CoCondenser+BM25) 7B hard 41.9 99.3 72.9 74.4
@@
LLM-QL (Full+LoRA) No Static(CoCondenser+BM25) 7B hard 42.4 99.4 73.6 72.4
@@ 7. Further Analyses
In this section, we conduct a thorough analysis of LLM-QL (full parameter fine-tuning for QL learning) to clarify its advantages.